### v1

@@ Abstract
Bayesian filtering is a key tool in many problems that involve the online processing of data, including data assimilation, optimal control, nonlinear tracking and others. Unfortunately, the implementation of filters for nonlinear, possibly high-dimensional, dynamical systems is far from straightforward, as computational methods have to meet a delicate trade-off involving stability, accuracy and computational cost. In this paper we investigate the design, and theoretical features, of constrained Bayesian filters for state space models. The constraint on the filter is given by a sequence of compact subsets of the state space that determines the sources and targets of the Markov transition kernels in the dynamical model. Subject to such constraints, we provide sufficient conditions for filter stability and approximation error rates with respect to the original (unconstrained) Bayesian filter. Then, we look specifically into the implementation of constrained filters in a continuous-discrete setting where the state of the system is a continuous-time stochastic Itô process but data are collected sequentially over a time grid. We propose an implementation of the constraint that relies on a data-driven modification of the drift of the Itô process using barrier functions, and discuss the relation of this scheme with methods based on the Doob $h$-transform. Finally, we illustrate the theoretical results and the performance of the proposed methods in computer experiments for a partially-observed stochastic Lorenz 96 model.
@@ 1.1 Motivation and background
The exact implementation of Bayesian filters is rarely feasible beyond the Kalman filter for linear and Gaussian models [Kalman60, Kalman61, Anderson79, Ristic04]. In general, approximations are needed, which include Kalman-based methods [Anderson79, Evensen03, Julier04], recursive Monte Carlo algorithms [Gordon93, Doucet00, Evensen03, Djuric03], variational schemes [Smidl08, Frazier23] and combinations of them. All these methods suffer from limitations which are related to a complex trade-off involving stability, accuracy and computational cost. For example, Kalman-based methods become very inaccurate when the system of interest displays strong nonlinearities or is subject to heavy-tailed perturbations [Ristic04], and particle filters (PFs) [Gordon93, Kitagawa96, Doucet00, Djuric03] usually struggle when the observations are highly-informative or the dimension of the model dimension is high [Bengtsson08, Snyder08].
@@
In this work, we are specifically interested in continuous-discrete filtering problems, where the state of the system of interest is a continuous-time stochastic process $\mathtt{X}(t)$ characterized by an Itô stochastic differential equation (SDE), and the observations $Y_{1},Y_{2},\ldots,Y_{n},$... are collected sequentially at specific time instants $t_{1},t_{2},\ldots,t_{n},$... [Akyildiz24]. In this setting, any constraint on the system dynamics needs to be translated into a modification of the underlying SDE. Such modifications are not straightforward. Related methods include guided SDEs, where an additional drift term is constructed using Doob’s $h$-transform [Pieper-Sethmacher25, Pieper-Sethmacher25b, Pieper-Sethmacher25c, Eklund25]. This transform, denoted $h(x,t)$, is generally intractable, however, and it is common to resort to linearizations of the original SDE in order to approximate $h(x,t)$. The methods in [Pieper-Sethmacher25, Pieper-Sethmacher25b, Pieper-Sethmacher25c, Eklund25] follow this approach. Moreover, they are developed for infinite-dimensional models and they enable sampling from regions of high probability given the observed data. The authors of [Chopin23], on the other hand, avoid linearizations and propose to approximate $h(x,t)$ using neural networks, which requires abundance of data and an offline training procedure. In any case, these techniques are computationally costly and subject to approximation errors that may be difficult to quantify.
@@ 1.2 Contributions
In this paper we investigate the problem of filtering for Markov state-space models (SSMs) [Sarkka13] under constraints. In a Markov SSM, the state dynamics is described by a sequence of Markov transition kernels $K_{1},K_{2},\ldots,K_{n},\ldots$ and we impose a constraint which is characterized by a sequence of compact subsets ${\mathcal{C}}_{0},{\mathcal{C}}_{1},\ldots,{\mathcal{C}}_{n},\ldots$ of the state space $\mathcal{X}$}. These sets specify the sources and targets of the kernels $K_{n}$, hence they yield a (possibly different) truncation of the state space $\mathcal{X}$} for each $n$. In this setup, we investigate the stability, approximation errors and numerical implementation of the resulting constrained Bayesian filters.
@@
    We parametrize the constraint by an index $l=1,2,\ldots$, in such a way that ${\mathcal{C}}_{n}^{l}\to{\mathcal{X}}$, for every $n$, when $l\to\infty$. Intuitively, the constraints on the state space are relaxed as $l$ increase, and we recover the original (unconstrained) model when $l\to\infty$.
@@
    We assume the compact subsets ${\mathcal{C}}_{0}^{l},{\mathcal{C}}_{1}^{l},\ldots,{\mathcal{C}}_{n}^{l},\ldots$ are sufficiently connected, in the sense that the probability mass allocated to the set ${\mathcal{C}}_{n}^{l}$ by the kernel $K_{n}$ when the initial state is a point ${\mathcal{C}}_{n-1}^{l}$ is bounded away from zero by some constant $\varepsilon_{n}^{l}>0$ that depends on $l$.
Under these assumptions, we find approximation error rates for the constrained filters with respect to $\varepsilon_{n}^{l}$. Moreover, when the positive bounds in assumption (b) above hold uniformly over time, i.e., $\varepsilon_{n}^{l}\geq\varepsilon^{l}>0$ for all $n$, the approximation rates can also be made uniform over $n$. Remarkably, the latter analysis suggests practical guidelines for the selection of the compact sets ${\mathcal{C}}_{0}^{l},{\mathcal{C}}_{1}^{l},\ldots,{\mathcal{C}}_{n}^{l},$... as superlevel sets of the likelihood functions in the unconstrained SSM.
A natural application of the results above is to study the approximation of Bayesian filters by constrained PFs, i.e., PFs that run on the constrained SSM. Under standard regularity assumptions, we obtain time-uniform error rates for PFs that depend both on the number of particles in the Monte Carlo filter and the lower bounds $\varepsilon^{l}>0$ for the transition probabilities from $x\in{\mathcal{C}}_{n-1}^{l}$ to ${\mathcal{C}}_{n}^{l}$ via $K_{n}$. Moreover, we show how the problem of weight degeneracy [Bengtsson08, Snyder08] can be mitigated in the constrained PF when the compacts ${\mathcal{C}}_{0}^{l},{\mathcal{C}}_{1}^{l},\ldots,{\mathcal{C}}_{n}^{l},$... are superlevel sets of the likelihood functions in the unconstrained SSM.
In the continuous-discrete filtering setup, the observations are collected at times $t_{1},t_{2},\ldots,t_{n},$..., and the Markov kernels $K_{n}$ are determined by the SDE that describes the continuous-time state process $\mathtt{X}(t)$ for $t\in[t_{n-1},t_{n}]$. We propose a simple (albeit possibly coarse) approximation of the constrained Markov kernels using so-called barrier functions [Margellos11, Clark21]. These functions can be computed online and they are used to modify the drift of the SDE in order to impose the constraint on the state space. As a result, we obtain a class of guided SDEs that can be easily simulated. Based on this approach, we show some numerical results that illustrate the stability and accuracy of the proposed constrained filters on a stochastic Lorenz 96 model [Grudzien20]
@@ 1.3 Organization of the paper
In Section 2 we introduce notation and provide some background on filter stability and constrained SSMs. Section 3 is devoted to the analysis of constrained Bayesian filters, including stability, approximation rates and convergence of constrained PFs. We note that Section 3 is essentially theoretical. Readers interested only in methodological aspects may skip directly to Section 4, where we describe the continuous-discrete filtering setup, discuss Doob’s $h$-transform for guided SDEs, and then introduce the proposed barrier-constrained SDEs. In Section 5 we present numerical results for selected constrained Bayesian filters (using the barrier approach) and the stochastic Lorenz 96 model. Finally, Section 6 is devoted to the conclusions.
@@ 2.1 State space models and Bayesian filtering
Let us construct a sequence of $d_{x}$-dimensional random vectors $\{X_{n}\}_{n\geq 0}$ on the probability space $(\Omega,{\mathcal{F}},\mathbb{P})$. The law of $X_{0}$ is denoted $\pi_{0}$ and the (random) dynamics of $X_{n}$ are determined by a sequence of Markov kernels
@@
where ${\mathcal{X}}\subseteq\mathbb{R}^{d_{x}}$ and ${\mathcal{B}}({\mathcal{X}})$ is the Borel $\sigma$-algebra of subsets of $\mathcal{X}$}. We also construct a sequence of uniformly bounded potential functions
@@
If $K=\{K_{n}\}_{n\geq 1}$ is the sequence of Markov kernels and $g=\{g_{n}\}_{n\geq 1}$ is the sequence of bounded potentials, then we can succinctly denote the system of interest as ${\mathcal{S}}:=(\pi_{0},K,g)$ and we refer to $\mathcal{S}$} as a Markov state space model (SSM).
Let ${\mathcal{P}}({\mathcal{X}})$ denote the family of probability measures on $({\mathcal{X}},{\mathcal{B}}({\mathcal{X}}))$ and, for any real function $f:{\mathcal{X}}\mapsto\mathbb{R}$ and $\mu\in{\mathcal{P}}({\mathcal{X}})$, let $\mu(f)$ be shorthand for $\int_{\mathcal{X}}f{\sf d}\mu$. We can define the sequence of operators $\Phi_{n}:{\mathcal{P}}({\mathcal{X}})\mapsto{\mathcal{P}}({\mathcal{X}}),$
(3) $\Phi_{n}(\mu)(f):=\frac{K_{n}\mu(fg_{n})}{K_{n}\mu(g_{n})}$
where $K_{n}\mu({\sf d}z):=\int_{{\mathcal{X}}}K_{n}(x,{\sf d}z)\mu({\sf d}x)$. If we let $\pi_{0}$ be the probability law of the initial state $X_{0}$, then the operators $\{\Phi_{n}\}_{n\geq 1}$ yield the sequence of probability measures
@@
In Bayesian filtering applications, the state sequence $\{X_{n}\}_{n\geq 0}$ is related to an observation sequence $\{Y_{n}\}_{n\geq 1}$, where $Y_{n}\in{\mathcal{Y}}\subseteq\mathbb{R}^{d_{y}}$, and $\mathcal{Y}$} is a $d_{y}$-dimensional observation space. In this setting, the potential $g_{n}(x)$ is the likelihood of the state $X_{n}=x_{n}$ given the observation $Y_{n}=y_{n}$, and $\pi_{n}$ is the conditional probability law of $X_{n}$ given $Y_{1}=y_{1},\ldots,Y_{n}=y_{n}$, often referred to as the optimal (Bayesian) filter at time $n$. Also, the measure
@@
Hereafter, we refer to $\Phi_{n}$ as a prediction-update (PU) operator, as it comprises the one-step-ahead prediction of (5) and the Bayesian update of (6). We also employ the term likelihood for $g_{n}$, rather than potential, in order to emphasize the relationship with the observations in the filtering problem.
@@ 2.2 A topology on the set of state space models
Let $\mathbb{S}$ be the set of SSMs of the form ${\mathcal{S}}=(\pi_{0},K,g)$ for which the sequence $\{\pi_{n}\}_{n\geq 1}$ is well defined. In particular, we assume that $K_{n}\pi_{n-1}(g_{n})>0$ for all $n\geq 1$, in order to guarantee that $\Phi_{n}(\pi_{n-1})=\pi_{n}\in{\mathcal{P}}({\mathcal{X}}_{n})$ is well defined.
@@
where the supremum is taken over all measurable subsets of $\mathcal{X}$}, denote the total variation norm of the difference $\alpha-\beta$. Following [Crisan20], we impose a topology $\mathbb{D}$ on the set of SSMs $\mathbb{S}$. To be specific, take some ${\mathcal{S}}\in\mathbb{S}$ and let ${\mathcal{S}}^{l}=(\pi_{0}^{l},K^{l},g^{l})$, $l\geq 1$, be a sequence is $\mathbb{S}$, with $K^{l}=\{K_{n}^{l}\}_{n\geq 1}$ and $g^{l}=\{g_{n}^{l}\}_{l\geq 1}$. We say that ${\mathcal{S}}^{l}$ converges to $\mathcal{S}$}, and denote
$\lim_{l\to\infty}{\mathcal{S}}^{l}={\mathcal{S}},$
when
@@
$\lim_{l\to\infty}\left\|K_{n}(x,\cdot)-K_{n}(x,\cdot)\right\|_{\rm tv}$ $=$ $0~~\text{for every $n\geq 1$, and}$
@@
Let $\{\Phi_{n}^{l}\}_{n\geq 1}$ denote the PU operators generated by model ${\mathcal{S}}^{l}$, which in turn yield the sequence of measures $\pi_{n}^{l}=\Phi_{0:n}(\pi_{0}^{l})$. The topology $\mathbb{D}$ has the property that convergence of the sequence of models ${\mathcal{S}}^{l}\to{\mathcal{S}}$ implies convergence of the probability laws $\pi_{n}^{l}\to\pi_{n}$ in total variation. We have the following lemma taken from [Crisan20].
@@
Let ${\mathcal{S}}^{l}=(\pi_{0}^{l},K^{l},g^{l})$, $l\geq 0$, and ${\mathcal{S}}=(\pi_{0},K,g)$ be elements of $\mathbb{S}$ with corresponding PU operators $\Phi_{n}^{l}$ and $\Phi_{n}$, $n\geq 1$, respectively. If $\lim_{l\to\infty}{\mathcal{S}}^{l}={\mathcal{S}}$, then $\lim_{l\to\infty}\left\|\Phi_{0:n}^{l}(\pi_{0}^{l})-\Phi_{0:n}(\pi_{0})\right\|_{\rm tv}=0$.

### v2

@@ Abstract
Bayesian filtering is a key tool in many problems that involve the online processing of data, including data assimilation, optimal control, nonlinear tracking and others. Unfortunately, the implementation of filters for nonlinear, possibly high-dimensional, dynamical systems is far from straightforward, as computational methods have to meet a delicate trade-off involving stability, accuracy and computational cost. In this paper we investigate the design, and theoretical features, of constrained Bayesian filters for state space models. The constraint on the filter is given by a sequence of compact subsets of the state space which determine the sources and targets of the Markov transition kernels in the dynamical model. Subject to such constraints, we provide sufficient conditions for filter stability, as well as approximation error rates with respect to the original (unconstrained) Bayesian filter. Then, we look specifically into the implementation of constrained filters in a continuous-discrete setting where the state of the system is a continuous-time stochastic Itô process but data are collected sequentially over a time grid. We propose an implementation of the constraint that relies on a data-driven modification of the drift of the Itô process using barrier functions, and discuss the relation of this scheme with methods based on the Doob $h$-transform. Finally, we illustrate the theoretical results and the performance of the proposed methods in computer experiments for a partially-observed stochastic Lorenz 96 model.
@@ 1.1 Motivation and background
The exact implementation of Bayesian filters is rarely feasible beyond the Kalman filter for linear and Gaussian models [Kalman60, Kalman61, Anderson79, Ristic04]. In general, approximations are needed, which include Kalman-based methods [Anderson79, Evensen03, Julier04], recursive Monte Carlo algorithms [Gordon93, Doucet00, Evensen03, Djuric03], variational schemes [Smidl08, Frazier23] and combinations of them. All these methods suffer from limitations which are related to a complex trade-off involving stability, accuracy and computational cost. For example, Kalman-based methods become very inaccurate when the system of interest displays strong nonlinearities or is subject to heavy-tailed perturbations [Ristic04], and particle filters (PFs) [Gordon93, Kitagawa96, Doucet00, Djuric03] usually struggle when the observations are highly-informative or the model dimension is high [Bengtsson08, Snyder08].
@@
In this work, we are specifically interested in continuous-discrete filtering problems, where the state of the system of interest is a continuous-time stochastic process $\mathtt{X}(t)$ characterized by an Itô stochastic differential equation (SDE), and the observations $Y_{1},Y_{2},\ldots,Y_{n},\ldots$ are collected sequentially at specific time instants $t_{1},t_{2},\ldots,t_{n},\ldots$ [Akyildiz24]. In this setting, any constraint on the system dynamics needs to be translated into a modification of the underlying SDE. Such modifications are not straightforward. Related methods include guided SDEs, where an additional drift term is constructed using Doob’s $h$-transform [Chopin23, Pieper-Sethmacher25, Pieper-Sethmacher25b, Pieper-Sethmacher25c, Eklund25]. This transform, denoted $h(x,t)$, is generally intractable, however, and it is common to resort to linearizations of the original SDE to approximate $h(x,t)$. The methods in [Pieper-Sethmacher25, Pieper-Sethmacher25b, Pieper-Sethmacher25c, Eklund25] follow this approach. Moreover, they are developed for infinite-dimensional models and they enable sampling from regions of high probability given the observed data. The authors of [Chopin23], on the other hand, avoid linearizations and propose to approximate $h(x,t)$ using neural networks, which requires abundance of data and an offline training procedure. In any case, these techniques are computationally costly and remain subject to approximation errors.
@@ 1.2 Contributions
We investigate the problem of filtering for Markov state-space models (SSMs) [Sarkka13] under constraints. In a Markov SSM, the state dynamics are described by a sequence of Markov transition kernels $K_{1},K_{2},\ldots,K_{n},\ldots$ and we impose a constraint which is characterized by a sequence of compact subsets ${\mathcal{C}}_{0},{\mathcal{C}}_{1},\ldots,{\mathcal{C}}_{n},\ldots$ of the state space $\mathcal{X}$}. These sets specify the sources and targets of the kernels $K_{n}$, hence they yield a specific truncation of the state space $\mathcal{X}$} for each $n$. In this setup, we look into the stability, approximation errors and numerical implementation of the resulting constrained Bayesian filters.
@@
    We parametrize the constraint by an index $l=1,2,\ldots$, in such a way that $\lim_{l\to\infty}{\mathcal{C}}_{n}^{l}={\mathcal{X}}$ for every $n$. Intuitively, this means that the constraints on the state space are relaxed as $l$ increases, and we recover the original (unconstrained) model when $l\to\infty$.
@@
    We assume that the compact subsets ${\mathcal{C}}_{0}^{l},{\mathcal{C}}_{1}^{l},\ldots,{\mathcal{C}}_{n}^{l},\ldots$ are sufficiently connected, in the sense that the probability mass $K_{n}(x,{\mathcal{C}}_{n}^{l})$, allocated to ${\mathcal{C}}_{n}^{l}$ by the kernel $K_{n}$ with initial point $x^{\prime}\in{\mathcal{C}}_{n-1}^{l}$, is bounded away from zero by some constant $\varepsilon_{n}^{l}>0$ that depends on $l$.
Under these assumptions, we find approximation error rates for the constrained filters with respect to $\varepsilon_{n}^{l}$. Moreover, when the positive bounds in assumption (b) above hold uniformly over time, i.e., $\varepsilon_{n}^{l}\geq\varepsilon^{l}>0$ for all $n$, the approximation rates can also be made uniform over $n$. Remarkably, this analysis suggests practical guidelines for the selection of the compact sets ${\mathcal{C}}_{0}^{l},{\mathcal{C}}_{1}^{l},\ldots,{\mathcal{C}}_{n}^{l},\ldots$ as superlevel sets of the likelihood functions in the unconstrained SSM.
In the continuous-discrete filtering setup, the observations are collected at times $t_{1},t_{2},\ldots,t_{n},\ldots$, and the Markov kernels $K_{n}$ are determined by the SDE that characterizes the continuous-time state process $\mathtt{X}(t)$ for $t\in[t_{n-1},t_{n}]$. We propose a simple (albeit possibly coarse) approximation of the constrained Markov kernels using barrier functions [Margellos11, Clark21]. These functions can be efficiently computed online and they are used to modify the drift of the SDE to impose the constraint on the state space. As a result, we obtain a class of modified SDEs that can be easily simulated and built into sampling-based algorithms such as PFs and ensemble Kalman filters (EnKFs). We show some numerical results that illustrate the stability and accuracy of the proposed constrained filters on a stochastic Lorenz 96 model [Grudzien20]
@@ 1.3 Organization of the paper
In Section 2 we introduce notation and provide some background on filter stability and constrained SSMs. Section 3 is devoted to the analysis of constrained Bayesian filters, including their stability and approximation rates. We note that Section 3 is essentially theoretical. Readers interested only in methodological aspects may skip directly to Section 4, where we describe the continuous-discrete filtering setup, discuss Doob’s $h$-transform for guided SDEs, and then introduce the proposed barrier-constrained SDEs. In Section 5 we present numerical results for selected constrained Bayesian filters (using the barrier approach) and the stochastic Lorenz 96 model. Finally, Section 6 is devoted to the conclusions.
@@ 2.1 State space models and Bayesian filtering
Let us construct a sequence of $d_{x}$-dimensional real random vectors $\{X_{n}\}_{n\geq 0}$ on the probability space $(\Omega,{\mathcal{F}},\mathbb{P})$. The law of $X_{0}$ is denoted $\pi_{0}$ and the (random) dynamics of $X_{n}$ are determined by a sequence of Markov kernels
@@
where ${\mathcal{X}}\subseteq\mathbb{R}^{d_{x}}$ is the state space and ${\mathcal{B}}({\mathcal{X}})$ is the Borel $\sigma$-algebra of subsets of $\mathcal{X}$}. We also construct a sequence of uniformly bounded potential functions
@@
If $K=\{K_{n}\}_{n\geq 1}$ is the sequence of Markov kernels and $g=\{g_{n}\}_{n\geq 1}$ is the sequence of uniformly bounded potentials, then we denote the system of interest as ${\mathcal{S}}:=(\pi_{0},K,g)$ and we refer to $\mathcal{S}$} as a Markov state space model (SSM).
Let ${\mathcal{P}}({\mathcal{X}})$ denote the family of probability measures on $({\mathcal{X}},{\mathcal{B}}({\mathcal{X}}))$ and, for any real function $f:{\mathcal{X}}\mapsto\mathbb{R}$ and $\mu\in{\mathcal{P}}({\mathcal{X}})$, let $\mu(f)$ be shorthand for $\int_{\mathcal{X}}f{\sf d}\mu$. We introduce the sequence of operators $\Phi_{n}:{\mathcal{P}}({\mathcal{X}})\mapsto{\mathcal{P}}({\mathcal{X}})$ such that
(3) $\Phi_{n}(\mu)(f):=\frac{K_{n}\mu(fg_{n})}{K_{n}\mu(g_{n})},$
where $K_{n}\mu({\sf d}z):=\int_{{\mathcal{X}}}K_{n}(x,{\sf d}z)\mu({\sf d}x)$. If we let $\pi_{0}$ be the probability law of the initial state $X_{0}$, then the operators $\{\Phi_{n}\}_{n\geq 1}$ yield the sequence of probability distributions
@@
In Bayesian filtering applications, the state sequence $\{X_{n}\}_{n\geq 0}$ is related to an observation sequence $\{Y_{n}\}_{n\geq 1}$, where $Y_{n}\in{\mathcal{Y}}\subseteq\mathbb{R}^{d_{y}}$, and $\mathcal{Y}$} is the $d_{y}$-dimensional observation space. In this setting, the potential $g_{n}(x)$ is the likelihood of the state $X_{n}=x$ given the observation $Y_{n}=y_{n}$, and $\pi_{n}$ is the conditional probability law of $X_{n}$ given $Y_{1}=y_{1},\ldots,Y_{n}=y_{n}$, often referred to as the optimal (Bayesian) filter at time $n$. Also, the measure
@@
Hereafter, we refer to $\Phi_{n}$ as the prediction-update (PU) operator, as it comprises the one-step-ahead prediction of (5) and the Bayesian update of (6). We also employ the term likelihood for $g_{n}$, rather than potential, to emphasize the relationship with the observations in the filtering problem.
@@ 2.2 A topology on the set of state space models
Let $\mathbb{S}$ be the family of SSMs of the form ${\mathcal{S}}=(\pi_{0},K,g)$ for which the sequence $\{\pi_{n}\}_{n\geq 1}$ is well defined (in particular, we assume that $\xi_{n}(g_{n})>0$, for all $n\geq 1$, in Eq. (6)).
@@
where the supremum is taken over all measurable subsets of $\mathcal{X}$}, denote the total variation norm of the difference $\alpha-\beta$. Following [Crisan20], we impose a topology $\mathbb{D}$ on the set of SSMs $\mathbb{S}$. To be specific, take some ${\mathcal{S}}\in\mathbb{S}$ and let ${\mathcal{S}}^{l}=(\pi_{0}^{l},K^{l},g^{l})$, $l\geq 1$, be a sequence in $\mathbb{S}$, with $K^{l}=\{K_{n}^{l}\}_{n\geq 1}$ and $g^{l}=\{g_{n}^{l}\}_{l\geq 1}$. ${\mathcal{S}}^{l}$ converges to $\mathcal{S}$}, and we denote $\lim_{l\to\infty}{\mathcal{S}}^{l}={\mathcal{S}},$ when
@@
$\lim_{l\to\infty}\left\|K_{n}(x,\cdot)-K_{n}^{l}(x,\cdot)\right\|_{\rm tv}$ $=$ $0~~\text{for every $n\geq 1$, and}$
@@
Let $\{\Phi_{n}^{l}\}_{n\geq 1}$ denote the PU operators generated by model ${\mathcal{S}}^{l}$, which in turn yield the sequence of measures $\pi_{n}^{l}=\Phi_{0:n}(\pi_{0}^{l})$. The topology $\mathbb{D}$ has the property that convergence of the sequence of models ${\mathcal{S}}^{l}\to{\mathcal{S}}$ implies convergence of the probability laws $\pi_{n}^{l}\to\pi_{n}$ in total variation. Specifically, we have the following lemma taken from [Crisan20].
@@
Let ${\mathcal{S}}^{l}=(\pi_{0}^{l},K^{l},g^{l})$, $l\geq 0$, and ${\mathcal{S}}=(\pi_{0},K,g)$ be elements of $\mathbb{S}$ with corresponding PU operators $\Phi_{n}^{l}$ and $\Phi_{n}$, $n\geq 1$, respectively. If $\lim_{l\to\infty}{\mathcal{S}}^{l}={\mathcal{S}}$ in $\mathbb{D}$, then $\lim_{l\to\infty}\left\|\Phi_{0:n}^{l}(\pi_{0}^{l})-\Phi_{0:n}(\pi_{0})\right\|_{\rm tv}=0$.